Circular No. 6897

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1989af IN UGC 6433 Jean Mueller reports her discovery of an apparent supernova (red mag

∼ 16.5) on a plate taken on 1989 Jan. 15 by P. Moniot with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. The star is located at α = 11^{h}25^{m}31^{s}.46, δ = +38^{o}03′43′′.6 (equinox 2000.0), which is 6′′.3 west and 6′′.1 north of the center of UGC 6433 (measures by G. V. Williams, SAO). SN 1989af is also present (mag 17.5–18) on films taken on 1989 Mar. 30 with the 0.46-m Palomar Schmidt telescope by H. Holt and N. G. Thomas. There is no object at this position on original Palomar Sky Survey prints, on the Digital Sky Survey, or on a second Survey IIIa-J plate taken on 1990 Apr. 22 by C. Brewer and J. D. Mendenhall.

SAX J1808.4–3658 = XTE J1808–369 A. V. Filippenko and D. C. Leonard, University of California at Berke-

ley, report: “Keck spectra (range 554–685 nm) of the suspected optical counterpart (cf. IAUC 6885, 6886) of SAX J1808.4–3658 reveal a possible very weak, double-peaked Hα emission line of velocity (full width) 1000 km/s. Absorption lines characteristic of mid- to late-type stars are present throughout the spectrum.”

M. Gilfanov, M. Revnivtsev, and R. Sunyaev, Max-Planck-Institut für Astrophysik and Russian Space Research Institute, report: “The analysis of the public-domain data of RXTE target-of-opportunity observations of this 2.5-ms binary x-ray pulsar (IAUC 6876–6878, 6885, 6886) during Apr. 11–May 2 has shown that an abrupt decline of the source luminosity occurred between Apr. 25 and 29. Whereas the luminosity was decreasing at an approximately constant rate corresponding to an e-folding time of ∼ 10 days between Apr. 11 and 25, it has dropped by a factor of ∼ 4 between Apr. 25 and 29, and by another factor of ∼ 10 by May 2. The RXTE observations have revealed surprising stability of the source spectrum during this luminosity decrease by more than two orders of magnitude. The source spectrum (energy range 3–100 keV for Apr. 11–29 data; 3–25 keV for May 2 data) had a remarkable Crab-like power-law shape with a photon index of ∼ 2, while the luminosity dropped from ∼ 8 × 10^{36} to ∼ 6 × 10^{34} erg/s (assuming a distance of 4 kpc). Such behavior is surprising within a framework of conventional models of the accretion onto a weakly magnetized neutron star. The shape of the x-ray light curve suggests that the magnetospheric radius has reached the corotation radius at the luminosity level of a few × 10^{35} erg/s.”

1998 May 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green